Hello I found this equation on SAT question of the day January 22, 2015
If f(x) = x + ax , and 7/2 what is f(3/2)
I chose D which was wrong
I do not understand where the 9/2x came from when clicking the correct answer
Please help correct and educate me
Thank you

Found 3 solutions by richard1234, MathLover1, MathTherapy:
Answer by richard1234(7193)

x + (7/2)x = (9/2)x, right? (hint: 1 + (7/2) = 9/2)

Therefore f(3/2) = (9/2)(3/2) = 27/4.

Answer by MathLover1(20855)
Answer by MathTherapy(10810)

Hello I found this equation on SAT question of the day January 22, 2015
If f(x) = x + ax , and 7/2 what is f(3/2)
I chose D which was wrong
I do not understand where the 9/2x came from when clicking the correct answer
Please help correct and educate me
Thank you

Too Ambiguous!!